

********************************************************************
UARS WEEKLY STATUS REPORT - 15 APR 2005
********************************************************************
***********
SOLSTICE
***********
This week SOLSTICE continued to operate in good health gathering solar
data on all available orbital tracking periods. On average, we received
63% of each day's data amounting to about 752 minutes of solar
observing time per day.
***********
SUSIM
***********
SUSIM resumed operations on April 12. Since then, the experiment
has successfully gathered daily mid (1.1 nm) resolution, MgII, and
continuous monitoring scans for time periods with UARS telemetry or
tape recording.
***********
PEM
***********
HEPSA data files are version 2. Currently, HEPSA data between 1991274
and 2004359 has been reprocessed and archived at the GSFC DAAC. HEPSA data
between 2004360 and 2004366 are currently waiting archival at GSFC DAAC.
MEPS data files are version 1. Processed MEPS data files at the DAAC
archive are between 1991274 and 2004359. MEPS data between 2004360 and 2004366
are currently waiting archival at GSFC DAAC.
PEM continues to receive NMC data files from NOAA. NMC host machine
changed during mid-January which caused some NMC data transfer failures;
however, this situation was corrected and recovery completed with the
assistance of Roger Lin at NOAA. No problems have been encountered since
updates have occurred.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional electron data for
the October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality. Proton data from these storms are now located at
http://climatology.space.swri.edu/UARS_images_from_Rudy_(Apr._2004).html
Requests have been received for storm spectral data. Overview plots were
produced and these plots may be found on the web at
http://climatology.space.swri.edu/UARS_images_postedJune2004.html
for the storms of October 29 of 1991, November 8-9 of 1991, April 6-7 of 2000,
March 31 - April 1 of 2001, April 11-12 of 2001, November 6-7 of 2001, and
November 11-12 of 2001.
PEM has received a request for data from the super storm of November
2004. The Dst value showed sudden commencement early on November 7, 2004
reaching a Dst value of about -380 gamma early on November 8, 2004.
PEM was not active at that time and became active late on November 8, 2004
in day only mode. A second peak of about -280 gamma was observed on November
11, 2004. PEM became active in day-light only mode late on November 8. PEM
survey spectrograms for November 8-18 of 2004 may be found at:
http://climatology.space.swri.edu/UARS_images_fromRudy.html
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 02 April 2005 (Orbit 74203, GMT Day 092) through
08 April 2005 (Orbit 74308, GMT Day 098).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 094/1453 74243 ZEPS/NEPS DAY-ONLY
SOLSTICE 095/1554 74258 DAYTIME TO FULLTIME OPS
The beta angle ranged between the angles of 46.7 degrees to 64.4
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 65 degrees on 09 April 2005 (DOY 099).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.16 to 3.82 degrees C. Additional
battery performance data for each battery on 01 April 2005 and 08
April 2005 are provided in Appendix D.
The clock error ranged from 12.1 to -10.3 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 08
PEM = 01
SOLSTICE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 05
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Five (5) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT continued to operate the batteries at the VT-5
charge rate. All indications showed the batteries to be performing
well at this level, so on the 095/1554 event SOLSTICE was powered on
fulltime. This was done to test the batteries abilities to sustain a
nighttime load. SOLSTICE is the smallest load out of the instruments,
so it was tried first. The batteries continued to look good with
SOLSTICE on fulltime, so the next step will be to try SUSIM (the next
biggest load) next week. The SUSIM turn on is scheduled for DOY 101.
Also this week, the FOT continued its testing of the solar array
drive. Once again both sides of the drive were utilized in order to
characterize their performance. The drive was moved 25 degrees in
either direction, however the results from these test runs were
inconsistent. Only the rewind time on the B drive and the ability to
move the drive in the forward direction via switching back and forth
between drives, seem to be repeatable. All other tests show different
results at different times. Due to the complexity of the drive itself
and the number of failure scenarios, it may take some time to isolate
the problem or problems. The FOT will continue to analyze the data
collected so far and run more tests on both drives next week.
Lastly, UARS flew through the solar eclipse on 08 April 2005
successfully. The array was moved prior to the first orbit impacted
by the eclipse in order to maximize daylight, which in turn minimized
the effect of the eclipse.
MEETINGS
The FOT held a meeting with the NASA mission directors on 07 April
2005 to report status on the current battery situation, array testing
and the course of action for next week.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 April 2005 (DOY 119). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 08 April 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
74243 094/1453 05-111 PEM ZEPS/NEPS DAY-ONLY
74243 094/1453 05-112 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
74245 094/1855 05-113 SADDE B-SIDE TESTING
74258 095/1544 05-114 SOLSTICE DAYTIME TO FULLTIME OPS
74260 095/1814 05-115 SADDE TESTING (ORBITS 74260-61)
74271 096/1231 05-116 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
74275 096/1741 05-117 SADDE TESTING
74287 097/1407 05-118 SADDE TESTING (THRU ORBIT 74289)
74296 098/0352 05-119 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
74303 098/1447 05-120 SADDE SADDE TESTING THRU ORBIT 74305
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 08 April 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 08 April 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
74243 094/1517 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
74271 096/1233 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
74296 098/0400 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 08 April 2005
01 April 2005 - GMT Day 091 Beta = 43.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.68/-4.94 +00.0/+00.0 06.9 0.00 0.0
2 291.2/ 89.6 +7.09/+5.79 +26.8/-14.4 25.8 1.07 15.4
3 78.4/-700.0 +2.60/+1.97 +20.8/-12.0 25.8 1.02 13.2
08 April 2005 - GMT Day 098 Beta = 64.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -4.31/-4.94 +00.0/+00.0 06.9 0.00 0.0
2 302.4/ 100.8 +5.79/+4.50 +19.6/-12.8 26.1 1.11 14.2
3 78.4/-403.2 +1.97/+1.34 +16.8/-12.0 26.1 1.06 13.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 08 April 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT 74215 092/1819 30832 Generic Late Acquisition
05-024 FOT 74247 094/2214 39145 No Acquisition
05-023 FOT 74252 095/0511 30832 Generic Late Acquisition
05-023 FOT 74300 098/1001 30832 Generic Late Acquisition
05-023 FOT 74301 098/1144 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-012 FOT See List Mar 05 30832 Generic Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-023 FOT See List Apr 05 30832 Generic Late Acquisition
05-024 FOT 74247 094/2214 39145 No Acquisition